Find the side lengths of each triangle.
It is an equalaterial triangle.
the sides are 2n+3.75 = 7n = 10n-2.25.
wat is the answer

Answer by Alan3354(21583)

2n+3.75 = 7n = 10n-2.25.
----------------
2n+3.75 = 7n
5n = 3.75
n = 0.75
---------
7n = 5.25
2n+3.75 = 5.25
10n-2.25 = 5.25
-------------------
Always be suspicious when there's more than one equal sign.